a block A weight of 320N resting on a rough horrizontal floor supports a block B 160 N. The two blcok connecting to a string passing arround a smooth pulley. Determine the horizontal force P that has to be applied on block A so as to just move it towards right.take miu=0.25 at all contact surface and also determine the tension T in the string respectively.
1
Expert's answer
2019-07-01T09:29:19-0400

According to Newton's second law, we can writer for a block A


"P=F_{friction} + T (1)"

wher P is the horizontal force, T is the tension in the string


Friction force is given by formula


"F_{friction}=kMg (2)"

The tension in the string is equal to


"T=mg (3)"

We put (2) and (3) in (1):

"P=kMg + mg (4)"

In our case we have


M=32 kg, m=16 kg, k=0.25


Using (4) we get:


P=240 N


Using (3) we get:


T=160 N


Answer:

P=240 N

T=160 N